OpenStudy (anonymous):

Hey guys, i have a test in about an hour, was just wondering if someone could help explain anti-derivatives a little. Thanks!

OpenStudy (zarkon):

it is just going backwards from differentiation

OpenStudy (anonymous):

ha yeah, i understand the simple ones, (like the antiderivative of 3x^2) would be x^3) but for the more advanced ones, is there a certain way to solve? if that makes sense haha

OpenStudy (zarkon):

what is your definition of advanced?

OpenStudy (anonymous):

ehh, i dont know just harder problems, im just not sure if there are certain steps to find the antiderivative or if i just kinda figure them out. Im looking over my notes now though and i think im starting to understand it a little more

OpenStudy (zarkon):

there are many techniques to finding antiderivatives...you need to be more specific on what you need help on

OpenStudy (zarkon):

probably the fist main rule you learned was antiderivatives of function of the form \[x^{n}\] its antiderivative is \[\frac{x^{n+1}}{n+1}\] provided that \[n\ne -1\]

OpenStudy (anonymous):

ok...well maybe for something like, 4+x^2 + 5x^3.....would it just be 4x + 1/3(x)^3 +5/4(x)^4 + (somenumber) ?

OpenStudy (anonymous):

i think that is what i was looking for! haha thanks!

OpenStudy (zarkon):

yes...use the letter c for the 'some number'

OpenStudy (anonymous):

okay

OpenStudy (zarkon):

\[4x+\frac{1}{3}x^3+\frac{5}{4}x^4+c\]

OpenStudy (anonymous):

awesome, thanks man appreciate it!

OpenStudy (zarkon):

no problem